OpenStudy (anonymous):

Simplify the following expression 3y-2x+z-2x+3(x+y)

OpenStudy (anonymous):

3y−2x+z−2x+3(x+y) Distribute: =3y+−2x+z+−2x+(3)(x)+(3)(y) =3y+−2x+z+−2x+3x+3y Combine Like Terms: =3y+−2x+z+−2x+3x+3y =(−2x+−2x+3x)+(3y+3y)+(z) =−x+6y+z Answer: =−x+6y+z

OpenStudy (ciarán95):

So, we need to look for common terms here within the expression. These are: x, y and z We'll group these together and tidy them up at the end. The first question we need to ask ourselves is...can we break down or simplify the expression? 3y - 2x + z - 2x + 3(x + y) We can multiply out the bracket at the end of the expression: 3(x + y). This will give us: 3y - 2x + z - 2x + 3x + 3y We can't simplify this any more, so all that's left to do is tidy it up by adding the common terms we mentioned at the start. To make this easier, we can rearrange the expression so that all of the common terms are beside one another. -2x - 2x + 3x + 3y + 3y + z We only have one term with 'z' in it, but we can add the 'x' and 'y' terms together: (-2 -2 + 3) + (3 + 3)y + z giving us: -x + 6y + z at the end!